Find two z values so that 38% of the middle area is bounded by them. Use a graphing calculator and round the answer to four decimal places, separated by a comma if needed.
Answer #1

P(-z < Z < z) = 0.38

P(Z < z) - P(Z < -z) = 0.38

P(Z < z) - ( 1 - P(Z < z) ) = 0.38

P(Z < z) - 1 + P(Z < z) = 0.38

2 P(Z < z) = 1.38

P(Z < z) = 0.69

From Z table, z-score for probability of 0.69 is 0.4959

z value = -0.4959 , 0.4959